A new plane cruises at a speed of 400 m.p.h. If it makes a trip in 7
hours with a strong wind and the return trip against the same wind in 9
hours, how fast was the wind?

Answer by Boreal(15235)

speed=400
wind=w
(400+w)7=7w+2800
(400-w)9=-9w+3600
those are equal, since they are the same distance
so 7w+2800=-9w+3600
16w=800
w=50 mph
450 mph with the wind*7 hours is 3150 miles
350 mph against the wind*9 hours is 3150 miles